Retailing Distribution Center Location Problem A major retail firm is considering a major change in its distribution system. It is considering relocating its distribution centers such that all retail stores can be replenished within 24 hours of a request for merchandise. The table below is a matrix of cover coefficients for 10 stores and 10 potential sites for distribution centers. If store i can be replenished within 24 hours from site j, then a_ij=1; otherwise a_ij-0. The cost of installation of the distributions centers vary from zone to zone and are given in the DC cost row. The firm wants to find the optimal location of the distribution centers.
Answer #1

1) Mathematical model is following

Decision variable: Xj = 1, if distribution center is located at site j, otherwise Xj = 0

Min 193X1+721X2+822X3+156X4+607X5+593X6+246X7+385X8+309X9+867X10

s.t.

X1+X2+X8+X9+X10 >= 1 (coverage for store 1)

X1+X2+X3 >= 1    (coverage for store 2)

X1+X2+X4+X5+X8+X9 >= 1 (coverage for store 3)

X3+X5+X6+X7+X9+X10 >= 1 (coverage for store 4)

X3+X4+X6+X7+X8+X9 >= 1 (coverage for store 5)

X4+X5+X6+X7+X10 >= 1   (coverage for store 6)

X2+X4+X6+X8+X9 >= 1    (coverage for store 7)

X1+X2+X3+X5+X8+X9+X10 >= 1     (coverage for store 8)

X1+X3+X4+X5+X7+X10 >= 1   (coverage for store 9)

X3+X6+X7+X8+X9+X10 >= 1   (coverage for store 10)

Xj \in {0,1}

2)

Decision variable: Xj = 1, if distribution center is located at site j, otherwise Xj = 0

Yi = 1 if store i can be replenished by at least one of the distribution centers, otherwise Yi = 0

Max Y1+Y2+Y3+Y4+Y5+Y6+Y7+Y8+Y8+Y9+Y10

s.t.

193X1+721X2+822X3+156X4+607X5+593X6+246X7+385X8+309X9+867X10 <= 300

X1+X2+X8+X9+X10 - Y1 >= 0 (coverage for store 1)

X1+X2+X3 - Y2 >= 0   (coverage for store 2)

X1+X2+X4+X5+X8+X9 - Y3 >= 0   (coverage for store 3)

X3+X5+X6+X7+X9+X10 - Y4 >= 0   (coverage for store 4)

X3+X4+X6+X7+X8+X9 - Y5 >= 0   (coverage for store 5)

X4+X5+X6+X7+X10 - Y6 >= 0   (coverage for store 6)

X2+X4+X6+X8+X9 - Y7 >= 0   (coverage for store 7)

X1+X2+X3+X5+X8+X9+X10 - Y8 >= 0 (coverage for store 8)

X1+X3+X4+X5+X7+X10 - Y9 >= 0 (coverage for store 9)

X3+X6+X7+X8+X9+X10 - Y10 >= 0 (coverage for store 10)

Xj, Yi  \in {0,1}

3)

Decision variable: Xj = 1, if distribution center is located at site j, otherwise Xj = 0

Yi = 1 if store i can be replenished by at least one of the distribution centers, otherwise Yi = 0

Min 193X1+721X2+822X3+156X4+607X5+593X6+246X7+385X8+309X9+867X10

s.t.

Y1+Y2+Y3+Y4+Y5+Y6+Y7+Y8+Y8+Y9+Y10 >= 10*0.85

X1+X2+X8+X9+X10 - Y1 >= 0 (coverage for store 1)

X1+X2+X3 - Y2 >= 0   (coverage for store 2)

X1+X2+X4+X5+X8+X9 - Y3 >= 0   (coverage for store 3)

X3+X5+X6+X7+X9+X10 - Y4 >= 0   (coverage for store 4)

X3+X4+X6+X7+X8+X9 - Y5 >= 0   (coverage for store 5)

X4+X5+X6+X7+X10 - Y6 >= 0   (coverage for store 6)

X2+X4+X6+X8+X9 - Y7 >= 0   (coverage for store 7)

X1+X2+X3+X5+X8+X9+X10 - Y8 >= 0 (coverage for store 8)

X1+X3+X4+X5+X7+X10 - Y9 >= 0 (coverage for store 9)

X3+X6+X7+X8+X9+X10 - Y10 >= 0 (coverage for store 10)

Xj, Yi  \in {0,1}

4)

Solution of model 1 using LINDO follows:

Lingo 17.0-Solution Report . Lingo! File Edit Solver Window Help 23Solution Report Lingol Lingo Model-Lingo Min- 193X1+721X

X1, X4, X7 = 1

Total cost = 595

____________________

Solution of model 2 using LINDO follows:

Lingo 17.0 Solution Report Lingol File Edit Solver Window Help Solution Report Lingo1 Lingo Model Lin Max = Y1+Y2+73+74+75+Y6Solution Report- Lingo1 Variable Yl Y2 Y3 Y4 YS Y6 Y7 Y8 Y9 Y10 X1 X2 X3 X4 X5 X6 X7 Value 1.000000 1.000000 1.000000 0.00000

X1 = 1,

Stores 1,2,3,8,9 are covered.

____________________

Solution of model 3 using LINDO follows:

Lingo 17.0 Solution Report Lingol File Edit Solver Window Help le- EZ Solution Report . Lingo! Lingo Model Lin Min 193 X1+721

Solution Report- Lingo1 Variable X1 X2 X3 X4 X5 X6 X7 X10 Yl Y2 Y3 Y4 YS Y6 17 Y8 Y9 Y10 value 1.000000 0.000000 0.000000 0.0

X1, X7 = 1

Total cost = 493

Except store 7, all stores are covered.